OpenStudy (anonymous):

Given cost and revenue functions, find the profit function. P(x) = R(x)-C(x)

OpenStudy (anonymous):

\[\huge R(x)= \frac{ 575 }{ 2\sqrt{x}}-3 \]

OpenStudy (anonymous):

\[\huge C(x)= 2000+150x-0.002x^2\]

OpenStudy (mertsj):

What seems to be the problem?

OpenStudy (anonymous):

@Mertsj i cant seem to subtract them

OpenStudy (mertsj):

That's because there are no like terms so why don't we just write: \[P(x)=\frac{575\sqrt{x}}{2x}+0.002x^2-150x-2003\]

OpenStudy (anonymous):



OpenStudy (anonymous):

@mertsj thats what my book says :S

OpenStudy (mertsj):

If that is the right answer, there is an error in the original posting.

OpenStudy (anonymous):

No there is no error i have checked the original question @mertsj

OpenStudy (mertsj):

Take a look at that.

OpenStudy (mertsj):

@burhan101